OpenStudy (richgall____):

Please Help Me With 2 questions :(

OpenStudy (xguardians):

How can I be of service :D?

OpenStudy (richgall____):

A pair of equations is shown below: y = 3x – 5 y = 6x – 8 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer.

OpenStudy (richgall____):

@xGuardians

OpenStudy (xguardians):

Okay, I prefer using substitution in this case, here are the steps. 1. Take a preferred equation and solve for x. y = 3x - 5 y + 6 = 3x (y+6)/3 = x Substitute that back into the other equation. y = 6(y+6/3) - 8 y = 2(y+6) - 8 y = 2y + 12 - 8 y = 2y + 4 -y = 4 y = -4 Subsitute this value into either equation to find x. -4 = 3x - 5 1 = 3x x = 1/3 To find where they intersect, simply set them equal to one another. 3x - 5 = 6x - 8 3x = 6x -3 3 = 3x x = 1 Substitute back into either equation to find y-value. y = 3(1) - 5 y = -2 At the point (1,-2), they intersect.